Quantitative Problem 1: Assume today is December 31, 2016. Barrington Industries expects that its 2017 after-tax operating income (EBIT(1-T)] will be $430 million and its 2017 depreciation expense will be $70 million. Barrington's 2017 gross capital expenditures are expected to be $100 million and the change in its net operating working capital for 2017 will be $20 million. The firm's free cash flow is expected to grow at a constant rate of 6.5% annually. Assume that its free cash flow occurs at the end of each year. The firm's weighted average cost of capital is 8.6%; the market value of the company's debt is $2.8 billion; and the company has 180 million shares of common stock outstanding. The firm has no preferred stock on its balance sheet and has no plans to use it for future capital budgeting projects. Using the free cash flow valuation model, what should be the company's stock price today (December 31, 2016)? Round your answer to the nearest cent. Do not round intermediate calculations. $ 30.34 per share

answer =43.08641975 is wrong according to Cengage.
Answer #1
FCFF = EBIT*(1-tax rate)+Depreciation- Capex- Chng. In Work. Cap.
FCFF = 430+70-100-20
FCFF = 380
firm or enterprise value = FCF in 1 year/(WACC - growth rate)
Firm/enterprise value = 380/ (0.086 - 0.065)
Firm/enterprise value = 18095.24
Enterprise value = Equity value+ MV of debt
18095.24 = Equity value+2800
Equity value = 15295.24
share price = equity value/number of shares
share price = 15295.24/180
share price = 84.97